A lot is 40 meters by 36 meters. Ahouse 27 meters by 9 meters is built on the lot.How much area is left over for a lawn?
Answer by tutorcecilia(2144)

Area = (length)(width) [Use the formula for the area of a rectangle]
Area (lot)=(40)(36)=1440 meters
Area (house)=(27)(9)=243
Area of lot - area of house = area of lawn
1440-243=1197 meters area for the lawn